# Generalized Face Liveness Detection via De-spoofing Face Generator

### v1

@@ 5 Conclusion
In this paper, we propose an Anomalous cue Guided Face Anti-Spoofing (AG-FAS) method. The anomalous cue is obtained via a De-spoofing Face Generator (DFG) trained only on real faces. The DFG model is capable of generating a “real” face version corresponding to the input. The residual of the generated “real” face image and the input image reflects the anomaly region of the input and serves as the anomalous cue for the downstream FAS task. As the DFG is trained on real faces with rich data diversity, the obtained anomalous cue is robust in different scenarios. Subsequently, we construct an Anomalous cue Guided FAS feature extraction Network (AG-Net) which introduces the embedded anomalous information into the feature extraction process at different layers via cross-attention modules. The AG-Net can benefit from the anomalous cue and obtain more generalized attack representations. We demonstrate the effectiveness of our method through extensive cross-domain FAS experiments with unseen scenarios and unknown presentation attacks.
@@
As this work solely relies on one additional real face dataset with 200K images to obtain the anomalous cue, one of our future directions will consider training the DFG on larger scale face datasets from the Internet, like Glint360K [56], to explore the possibility of further improving generalization.

### v2

@@ VI Conclusion
In this paper, we propose an Anomalous cue Guided Face Anti-Spoofing (AG-FAS) method, which can effectively leverage additional data of real faces for improving model generalization on cross-domain FAS tasks. We train a De-fake Face Generator (DFG) on a large-scale real face only dataset, enabling it to generate the “real” version of any input face. The residual of the generated “real” face and the input reflects the off-real region and serves as the anomalous cue for FAS. Specifically, we construct an Off-real Attention Network (OA-Net) that utilizes the anomalous cue to allocate its attention toward the off-real region in the input, achieving a more generalized FAS feature. We demonstrate the effectiveness of our method through both theoretical analysis and extensive cross-domain FAS experiments with unseen scenarios and unknown presentation attacks.
@@
As this work solely relies on two additional datasets with 130K real faces to obtain the anomalous cue, one of our future directions will consider training the DFG on larger scale face datasets from the Internet, like Glint360K [68], to explore the possibility of further improving generalization.